# MET-Bench: Multimodal Entity Tracking for Evaluating the Limitations of Vision-Language and Reasoning Models

### v1

@@ Abstract
Entity tracking is a fundamental challenge in natural language understanding, requiring models to maintain coherent representations of entities. Previous work has benchmarked entity tracking performance in purely text-based tasks. We introduce MET-Bench, a multimodal entity tracking benchmark designed to evaluate the ability of vision-language models to track entity states across modalities. Using two structured domains, Chess and the Shell Game, we assess how effectively current models integrate textual and image-based state updates. Our findings reveal a significant performance gap between text-based and image-based tracking and that this performance gap stems from deficits in visual reasoning rather than perception. We further show that explicit text-based reasoning strategies improve performance, yet substantial limitations remain, especially in long-horizon multimodal scenarios. Our results highlight the need for improved multimodal representations and reasoning techniques to bridge the gap between textual and visual entity tracking.
MET-Bench: Multimodal Entity Tracking for Evaluating
the Limitations of Vision-Language and Reasoning Models
Vanya Cohen  and Raymond Mooney Department of Computer Science The University of Texas at Austin {vanya,mooney}@utexas.edu
@@ 1 Introduction
Natural language understanding requires the ability to track and update information about entities as they evolve through text. From coreference resolution (Hobbs, 1978; Lappin and Leass, 1994) and discourse processing to narrative comprehension, computational linguistics has long grappled with the challenge of maintaining coherent entity representations across textual contexts (Bunescu and Paşca, 2006; Schank and Abelson, 1977).
@@
While significant progress has been made in tasks like coreference resolution and entity linking Liu et al. (2023); Papalampidi et al. (2022), the broader challenge of tracking entity states—understanding how entities change through sequences of actions or events—remains an open challenge (Fagnou et al., 2024; Kim and Schuster, 2023; Toshniwal et al., 2022). This limitation becomes particularly apparent in tasks requiring integration of information across multiple modalities, an increasingly important frontier in computational linguistics as language processing systems are asked to reason about content that combines text with other forms of communication like images and video.
@@
Our work examines this challenge through the lens of multimodal entity state tracking, where changes to entity states must be understood from both textual descriptions and visual observations. This setting provides a natural extension to classical NLP problems like discourse processing and situated language understanding, while also connecting to emerging research in multimodal dialogue and human-AI interaction. We focus specifically on scenarios where language models must reason about world-state changes described through a combination of text and images. Consider the task of understanding assembly instructions that combine written steps with supporting diagrams: while text might specify "Attach panel A to the base using the provided screws," accompanying images show the precise alignment and orientation. Accurate language understanding in such contexts requires maintaining a coherent mental model that integrates both linguistic and visual information about how entities’ states evolve.
(a) Chess State
(b) Chess Action
(c) Shell Game State
(d) Shell Game Action
Figure 1: Illustration of the two domains used in this work. (a) An example Chess board state. (b) The chess move (action) rendered as an image. (c) An initial shell game state with the blue ball under a shell. (d) Image action representing shells at positions one and three being swapped.
@@
To systematically evaluate models’ capabilities in this multimodal language understanding setting, we introduce two complementary benchmarks: multimodal Chess and the Shell Game. Through these domains, we assess how effectively current language models can track entity states when updates are conveyed through both text and images. Our analysis reveals substantial disparities in how models process text-based and image-based entity-state updates, highlighting fundamental limitations in their multimodal language understanding.
@@
While Chess and the Shell Game provide structured testbeds for evaluating entity tracking, real-world applications often involve more ambiguous and dynamic environments. However, by isolating state-tracking performance in controlled settings, we establish a clear baseline for assessing multimodal reasoning, one that provides a straightforward means of evaluation and can scale in difficulty with minimal changes.
We hypothesize that current language models struggle with multimodal entity tracking not due to low-level perceptual failures but because they lack representations (learned or otherwise engineered) for maintaining entity coherence across sequential visual observations. This suggests a fundamental limitation in how these models integrate and update state representations from different modalities.
We make the following contributions:
@@
  • Through probing experiments, we show that these limitations stem from higher-level reasoning challenges rather than low-level perception issues.
  • We evaluate various approaches to improving multimodal entity tracking, finding that techniques emphasizing explicit reasoning outperform traditional NLP methods like fine-tuning when generalizing to novel domains.
@@ 2 Background
We formulate the problem of multimodal entity tracking as a sequential state estimation task, where an agent must infer the final state of a system given an initial state and a series of observed actions. Formally, at each time step $t$, the environment is in state $\bm{S_{t}}$, and transitions occur according to an action sequence $\bm{A}=(\bm{a_{1}},\bm{a_{2}},\dots,\bm{a_{T}})$. The agent receives observations $\bm{A_{t}}$ corresponding to each action, which may be textual ($\bm{A_{t}^{\text{text}}}$) or visual ($\bm{A_{t}^{\text{image}}}$). The objective is to infer the final state,
$\bm{S_{T}}=f(\bm{S_{0}},\bm{A_{1}},\bm{A_{2}},\dots,\bm{A_{T}}),$
where $f$ is the function modeling entity state updates.
MET-Bench represents the initial and final states of each domain as text but evaluate the models’ ability to track entity state changes through images. This approach isolates the multimodal entity tracking challenge by ensuring that models begin and end with well-defined textual representations while processing state transitions visually. By doing so, we assess their capacity to maintain coherent entity representations across modalities while minimizing confounding errors from perceptual failures, which remain a known limitation of current vision-language models Sharma et al. (2024).
@@ 2.1 Chess Domain
Chess is a well-studied domain for testing entity tracking of deep learning models Toshniwal et al. (2022). The state $S_{t}$ represents an 8×8 board configuration expressed in Forsyth–Edwards Notation (FEN) notation, actions correspond to legal chess moves from real games, and action observations consist of either symbolic (UCI notation) or visual (board images) descriptions of moves. We likewise adopt chess as an entity tracking testbed where the task is to maintain a correct representation of the board state across a sequence of moves. This distinction foregrounds how well a model can integrate and track piece locations as they change over time in potentially complicated board configurations.
Utilizing real Chess games from the Millionbase dataset [^1] https://rebel13.nl/rebel13/rebel%2013.html used in Toshniwal et al. (2022), we generate sequences of states and actions (moves) using standard chess notation: Universal Chess Interface (UCI) for actions and FEN for board states.
  • Text-Encoded Moves: Each action is provided as a short UCI textual description (e.g., ‘‘e2e4’’ for moving a piece from e2 to e4).
  • Image-Encoded Moves: Each action is accompanied by a rendered image that serves as a visual representation of the move (see Fig. 1).
@@ 2.2 Shell Game Domain
The second domain in MET-Bench is the Shell Game, a classic demonstration of hidden-state tracking. A ball is placed under one of three cups (or shells), which are then swapped pairwise in succession. The goal is tracking which cup currently hides the ball as shells are swapped. The state $S_{t}$ tracks the hidden position of a ball under three shells, actions correspond to swaps between pairs of shells. Other works have explored shell-game-like domains with varying levels of added complexity (Li et al., 2021; Long et al., 2016; Kim and Schuster, 2023).
This domain has a simpler entity-state and action space than Chess. However, while many frontier models have been trained on UCI/FEN encoded chess games, the Shell Game is, to the best of our knowledge, not present in the training data of these models. There may however be analogous tasks in the pre-training data.
We simulate repeated Shell Game swaps to create a set of Shell Game trajectories. Swap actions are either:
  • Text-Encoded Swap: Denoted as ‘‘x swap y’’, where $\{x,y\}$ are in $\{1,2,3\}$.
  • Image-Encoded Swap: An image depicting the shells being swapped, with the ball visually hidden (see Fig. 1).
The ground truth entity state after the game finishes is a single number indicating the final shell position of the ball.
2.3 Models
We use MET-Bench to evaluate the limitations of frontier models, including vision-language models (VLMs) which accept images and text as input, and newer reasoning models like OpenAI’s o1 that are trained using reinforcement learning and utilize test-time search algorithms to improve their reasoning abilities on domains like mathematics and coding. A full list of models and their capabilities is shown in Appendix A, Table 5.
@@ 3 Methods
Now what is the final FEN? Only output the FEN.
Assistant rnbqkbnr/pppp1ppp/8/4p3/4P3/ 8/PPPP1PPP/RNBQKBNR w KQkq - 2 2
@@
We utilize the standard chat-based schema exposed by current frontier models that consists of interleaved user-provided and assistant (model) provided messages. Figures 2 shows the prompting strategy used for the Chess domain. Similarly, Appendix A, 6 shows the prompting strategy used for the Shell Game domain. Text actions are represented using simple notation on which the models have been trained, UCI for Chess and a simple domain-specific-language for Shell Game.
In the case of image-action input, the text actions are replaced with their image-rendered versions as Base64 encoded PNG images and a text description of how to interpret the image-actions is provided. Fig. 1 shows the image representations used for the Chess and Shell Game actions. These image representations were created through visual-prompt engineering to maximize the classification accuracy of actions depicted. Various common image representations were explored including arrows, bounding boxes, and symbolic markers. The image depiction of the game actions is explained to the language model every time images are provided using the prompts in Appendix A, Figures 7 & 8.
@@ 4 Experiments
Chess Text Image
@@
Claude 3.5 Sonnet 96.8 66.2
MiniMax-VL-01 86.3 65.8
Gemini-2.0-Flash 93.3 74.7
GPT-4o mini 68.3 60.6
GPT-4o 89.6 73.3
@@
Claude 3.5 Sonnet 99.6 77.9
MiniMax-VL-01 88.0 75.9
Gemini-2.0-Flash 96.6 78.7
GPT-4o mini 75.1 74.9
GPT-4o 94.9 77.7
@@
Claude 3.5 Sonnet 97.9 68.4
MiniMax-VL-01 62.7 58.0
Gemini-2.0-Flash 93.3 74.5
GPT-4o mini 72.0 69.6
GPT-4o 91.8 74.9
@@
o1-mini 65.1 -
o3-mini 99.6 -
o1 98.2 83.5
(a) Chess with ten moves in each sequence.
Shell Game Text Image
@@
Claude 3.5 Sonnet 34.4 36.2
MiniMax-VL-01 34.4 35.2
Gemini-2.0-Flash 30.0 33.4
GPT-4o mini 30.6 32.8
GPT-4o 36.0 32.4
@@
Claude 3.5 Sonnet 34.0 30.6
MiniMax-VL-01 36.4 32.0
Gemini-2.0-Flash 37.0 31.4
GPT-4o mini 34.4 31.2
GPT-4o 37.2 31.0
@@
Claude 3.5 Sonnet 97.4 94.2
MiniMax-VL-01 92.6 32.8
Gemini-2.0-Flash 76.8 33.8
GPT-4o mini 84.4 35.0
GPT-4o 99.8 84.2
@@
o1-mini 99.8 -
o3-mini 100.0 -
o1 100.0 92.8
(b) Shell Game with five swap moves in each sequence.
Table 1: Entity tracking accuracy in Chess and Shell Game for text and image actions. In the Few-Shot setting $N=5$ in-context examples are used. Methods and models which employ explicit reasoning perform best (chain-of-thought and reasoning models).
4.1 Tracking Entities with Text and Image Actions
We evaluate difference in accuracy when tracking images from text and image actions in the zero-shot, few-shot, and chain-of-thought settings. These results are presented in Section 5.1.
Zero-shot
Chess
In the Chess domain, we evaluate on a set of $100$ games selected at random from the test set, each with a sequence length of ten actions. The model must predict the FEN string for the final state. If the FEN string contains syntax errors such that it cannot be parsed, the accuracy for that instance is zero. We report the per-square accuracy of the predicted board, that is the ratio of correctly predicted pieces (or absence of a piece) to the total number of board tiles. The ‘Game Start’ baseline is the accuracy of predicting the initial board configuration. After only ten actions, most of the board configuration remains unchanged, so this is a strong baseline.
Shell Game
@@ 4.1 Tracking Entities with Text and Image Actions
We evaluate Shell Game using a set of $500$ games generated at random, each with a sequence length of five actions. The final state is a single number $n\in\{1,2,3\}$ that gives the position of the ball, and we measure the accuracy of predicting it. The naive baseline picks a position uniformly at random.
Few-Shot and Chain-of-Thought
In these settings, the evaluation and procedure remain largely unchanged from the zero-shot Chess. For the few-shot experiments, $N=5$ examples are selected at random from the training set and prepended to the test example. The few-shot examples have the same number of actions as the test-set examples. To evaluate the effect of chain-of-thought reasoning, we prompt the model to ‘think step by step before producing a final answer.’
4.2 Sequence-Length Variation
The sequence length of the actions is varied to quantify the effect of compounding errors on the models. These sequences range from zero to $100$ actions in both the image and text action modality. This serves to quantify the relative drop-off in performance among models in the zero-shot setting. These results are presented in Section 5.2.
4.3 Image-Action Classification
We perform an experiment to demonstrate that VLMs have the ability to accurately interpret the actions depicted in the image-action representations. Figures 7 & 8 in Appendix A show the prompting strategy used to get a VLM to predict (zero-shot) a text action from an image action. We perform this evaluation on a test set of $1000$ image depictions of actions for each game. The results are presented in Section 5.3.
4.4 Cascaded Inference
Using the text actions predicted from the images in the image-action classification task, we devise an ablation to test the effect of cascading (performing multimodal inference in two steps) in the zero-shot setting. Given a sequence of image-based observations $\bm{A_{t}^{\text{image}}}$, a vision-language model (VLM) first predicts the corresponding text-based action sequence $\hat{\bm{A}}^{\text{text}}=g(\bm{A}^{\text{image}})$, where $g$ maps images to text using the procedure in Section 4.3. The text actions are then used for zero-shot inference to estimate the final state as $\bm{S_{T}}=f(\bm{S_{0}},\hat{\bm{A}}^{\text{text}})$. This removes the need for the model to perform entity tracking directly in the image modality, isolating the effect of perceptual failures. These results are presented in Section 5.4.
4.5 Fine-Tuning Experiments
We fine-tune frontier OpenAI models using their fine-tuning API on a small training set. We evaluate these in the Chess and Shell Game domains using both image and text modalities. For Chess, we select a random set of $100$ training examples of sequence length ten. For Shell Game we use $20$ training examples of sequence length $3$. Results are available in Section A.2. Hyperparameters were determined through grid search and are available in Appendix A.2, Table 6.
4.6 Mixed-Modality Experiments
We perform an ablation where the image-action and text-action modalities are mixed. At each action step, one action modality is randomly selected with a probability varying from $100\%$ text actions to $100\%$ image actions. This serves to quantify the effect of forcing the model to reason over multiple input modalities simultaneously. These results are presented in Section A.1.
5 Results
5.1 Tracking in Text Outperforms Images
Chess
Table 1(a) reports the accuracy for text and image actions in the chess domain. Performance in the text modality is significantly better than in the image modality. In the zero-shot setting, the best-performing text model, Claude 3.5 Sonnet, achieves an impressive $96.8\%$ accuracy, while its image-based counterpart drops sharply to $66.2\%$. A similar trend holds across models, indicating that current models can perform entity tracking in text but fail to integrate visual updates as effectively.
Both in-context learning with few-shot prompting and chain-of-thought reasoning lead to performance improvements. In the image modality, the accuracy is not significantly greater than the naive baseline. Only the o1 reasoning model achieves an accuracy greater than $80\%$ in the image modality.
Shell Game
(a) Chess zero-shot accuracy with text actions.
(b) Chess zero-shot accuracy with image actions.
Figure 3: In the text action setting, the reasoning models, o1 and o3-mini, maintain the highest accuracy at longer sequence lengths. o1-mini and the other models begin to output invalid and inaccurate board representations. All models struggle to maintain accurate board representations in the image action setting, with o1 performing the best.
@@ 5.1 Tracking in Text Outperforms Images
Table 1(b) reports the accuracy for text and image actions for Shell Game. The results follow a pattern similar to that of Chess, with text-based tracking outperforming image-based tracking. However, unlike in Chess, the performance in the zero-shot setting is close to random. But the best model, o1, attains accuracies of $100.0\%$ and $92.8\%$ for the text and image modalities, respectively.

### v2

@@ Abstract
Entity state tracking is a necessary component of world modeling that requires maintaining coherent representations of entities over time. Previous work has benchmarked entity tracking performance in purely text-based tasks. We introduce MET-Bench, a multimodal entity tracking benchmark designed to evaluate vision-language models’ ability to track entity states across modalities. Using two structured domains, we assess how effectively current models integrate textual and image-based state updates. Our findings reveal a significant performance gap between text-based and image-based entity tracking. We empirically show this discrepancy primarily stems from deficits in visual reasoning rather than perception. We further show that explicit text-based reasoning strategies improve performance, yet limitations remain in long-horizon multimodal tasks. We develop a reinforcement learning method to improve performance on MET-Bench. Applying our method to open-source VLMs achieves competitive performance with advanced closed models. Our results highlight the need for improved multimodal representations and reasoning techniques to bridge the gap between textual and visual entity tracking.
Machine Learning, ICML
@@ 1 Introduction
Developing models that track and predict the latent state of the world has emerged as an important goal for artificial intelligence (worldmodels2024). One faculty of effective world models is entity state tracking: estimating the evolution of entities, their attributes and relations over time. Many problems require accurate entity tracking including robotic manipulation (yoneda2024statler; nvidia2025gr00tn1openfoundation), video–question‐answering (he2024mmworld), and computer-control agents (bishop2024latentstateestimationhelps; sager2025comprehensive). Success on these tasks requires maintaining a coherent estimate of how entity state changes from sequential, multimodal observations.
@@
Early progress on entity tracking emerged within text-only tasks. From coreference resolution (HOBBS1978311; Lappin and Leass, 1994) and discourse processing to narrative comprehension, computational linguistics has long grappled with the challenge of maintaining coherent entity representations across textual contexts (Bunescu and Paşca, 2006; Schank1977).
@@
While significant progress has been made in text-only tasks (liu2023brief; papalampidi2022towards), the broader challenge of tracking entity states—understanding how entities change through sequences of actions or events—remains an open challenge (Fagnou et al., 2024; Kim and Schuster, 2023; toshniwal2022chess). This limitation becomes particularly apparent in tasks requiring integration of information across multiple modalities, an increasingly important frontier in AI as systems are asked to reason about content that combines text with other modalities like images and video.
@@
Our work examines this challenge through the lens of multimodal entity state tracking, where changes to entity states must be understood from both textual descriptions and visual observations. This setting provides a natural extension to classical NLP problems while also connecting to emerging research in multimodal dialogue, robotics, and human-AI interaction. We focus specifically on scenarios where multimodal models must reason about world-state changes described through a combination of text and images.
Figure 1: In the multimodal entity tracking benchmark (MET-Bench), a vision-language model (VLM) predicts the final entity state from the initial state and actions which update the entity state. The initial entity state is provided as text and the actions are given as images or as text. The predicted final state is a text representation of the entity state.
@@
To systematically evaluate models’ capabilities in this multimodal setting, we introduce two complementary benchmarks: multimodal Chess and Shell Game. Through these domains, we assess how effectively current language models can track entity states when updates are conveyed through both text and images. Our analysis reveals substantial disparities in how models process text-based and image-based entity-state updates, highlighting fundamental limitations in their multimodal understanding.
@@
While Chess and the Shell Game provide structured testbeds for evaluating entity tracking, real-world applications often involve more ambiguous and dynamic environments. However, by isolating state-tracking performance in controlled settings, we establish a clear baseline for assessing multimodal reasoning, one that provides a straightforward means of evaluation where difficulty is easily controlled.
Overall, this paper makes the following contributions:
@@
  • Through experiments, we show that these limitations stem from higher-level reasoning challenges rather than simple perceptual issues.
  • We apply reinforcement learning to improve the reasoning of open-source VLMs for entity tracking, and achieve performance competitive with advanced closed models. However, we find only limited transfer learning between modalities.
@@ 2 Background
We formulate the problem of multimodal entity tracking as a sequential state estimation task, where an agent must infer the final state of a system given an initial state and a series of observed actions. Formally, at each time step $t$, the environment is in state $\mathbf{S}_{t}$, and transitions occur according to an action sequence $\mathbf{A}=(\mathbf{a}_{1},\mathbf{a}_{2},\dots,\mathbf{a}_{T})$. The agent receives $\mathbf{a}_{t}$ corresponding to each action, which may be textual $\mathbf{a}_{t}^{\text{text}}$ or visual $\mathbf{a}_{t}^{\text{image}}$. The objective is to infer the final state,
$\mathbf{S}_{T}=\operatorname{f}(\mathbf{S}_{0},\mathbf{a}_{1},\mathbf{a}_{2},\dots,\mathbf{a}_{T}),$
where $\operatorname{f}$ is the function modeling entity state updates.
MET-Bench represents the initial and final states of each domain as text but also evaluates models’ ability to track entity state changes through images. This approach isolates the multimodal entity tracking challenge by ensuring that models begin and end with well-defined textual representations while processing state transitions visually. This assesses the models’ capacity to maintain coherent entity representations across modalities while minimizing confounding errors from perceptual failures, which remain a limitation of current vision-language models (sharma2024vision).
@@ 2.1 Chess Domain
Chess is a well-studied domain for testing entity tracking of deep learning models (toshniwal2022chess). The state $S_{t}$ represents an 8×8 board configuration expressed in Forsyth–Edwards Notation (FEN) notation, actions correspond to legal chess moves from real games, and action observations consist of either symbolic (UCI notation) or visual (board images) descriptions of moves. We likewise adopt chess as an entity tracking testbed where the task is to maintain a correct representation of the board state across a sequence of moves.
Utilizing real games from Millionbase [^1] https://rebel13.nl/rebel13/rebel%2013.html (toshniwal2022chess), we generate sequences of states and actions (moves) using standard chess notation: Universal Chess Interface (UCI) for actions and FEN for board states. Text-Encoded Moves (each action is provided as a short UCI textual description, e.g., ‘‘e2e4’’ for moving a piece from e2 to e4); Image-Encoded Moves (each action is accompanied by a rendered image serving as a visual representation of the move; see Fig. 1). The visual action depictions were created to maximize perceptual accuracy as in Table 4.
@@ 2.2 Shell Game Domain
The second domain is Shell Game, a classic demonstration of hidden-state tracking. A ball is placed under one of three cups (or shells), which are then swapped pairwise in succession. The goal is tracking which cup currently hides the ball. The state $S_{t}$ tracks the hidden position of a ball under three shells, actions correspond to swaps between pairs of shells. Other works have explored shell-game-like domains with varying levels of added complexity (Li et al., 2021; Long et al., 2016; Kim and Schuster, 2023).
Shell Game has a simpler entity-state and action space than Chess. However, while many frontier models have been trained on UCI/FEN encoded chess games, the Shell Game is, to the best of our knowledge, not present in the training data of these models, though there may be analogous tasks in the pre-training data.
We simulate repeated Shell Game swaps to create a set of Shell Game trajectories. Swap actions are either: Text-Encoded Swaps (‘‘x swap y’’, where $\{x,y\}\subset\{1,2,3\}$); Image-Encoded Swaps (an image depicting the shells being swapped, with the ball not visible; see Fig. 1). The ground truth entity state after the game finishes is a single number indicating the final shell position of the ball.
2.3 Reinforcement Learning for Reasoning
To improve the entity state reasoning of open-source models, we utilize Group Relative Policy Optimization (GRPO) (shao2024deepseekmath), a policy-gradient RL method for training language models.
GRPO is suited to tasks with verifiable outcomes, such as entity tracking, where correctness can be determined programmatically. By using the task’s ground-truth accuracy as the reward signal, GRPO can directly optimize for entity tracking performance without requiring human preference data or learned reward models.
@@ 3 Methods
Now what is the final FEN? Output FINAL ANSWER: [FEN].
Assistant FINAL ANSWER: "rnbqkbnr/pppp1ppp/8/4p3/4P3 /8/PPPP1PPP/RNBQKBNR w KQkq - 0 2"
@@
We utilize the standard chat-based input schema exposed by current models that consists of interleaved user-provided and model-generated messages. Figure 2 shows the prompting strategy used for the Chess domain. Similarly, Appendix A details the prompting strategy used for the Shell Game domain.
In the case of image-action input, the text actions are replaced with their image-rendered versions as Base64 encoded PNG images and a text description of how to interpret the image-actions is provided. Fig. 1 shows the image representations used for the Chess and Shell Game actions. These image representations were created through visual-prompt engineering to maximize the classification accuracy of actions depicted. Various common image representations were explored including arrows, bounding boxes, and symbolic markers. The image depiction of the game actions is explained to the language model every time images are provided using the prompts in the Appendix.
3.1 GRPO Training for Entity Tracking
To investigate whether reinforcement learning can improve entity tracking in open-weight vision-language models, we apply GRPO to finetune models on MET-Bench tasks.
Reward Function Design
For entity tracking, we design reward functions that leverage the automatic verifiability of our benchmark tasks. For Chess, the reward is the fraction of correctly predicted board squares:
$R_{\text{chess}}(y,y^{*})=\frac{1}{64}\sum_{i=1}^{64}\mathbbm{1}[y_{i}=y^{*}_{i}],$
where $y$ is the predicted FEN board state and $y^{*}$ is the ground truth. For the Shell Game, the reward is binary based on correct ball position prediction:
$R_{\text{shell}}(y,y^{*})=\mathbbm{1}[y=y^{*}].$
We train Gemma3 4B and 27B (gemmateam2025gemma3technicalreport). Full training configurations are in Appendix A.
(a) Chess accuracy with text actions.
(b) Chess accuracy with image actions.
Figure 3: In the text action setting, the reasoning models Claude 3.7 Sonnet Thinking and Gemini 2.5 Pro, maintain the highest accuracy at longer action-sequence lengths. All models struggle to maintain accurate board representations in the image action setting, with Claude 3.7 Sonnet Thinking performing the best. Bars show 95% confidence intervals.
(a) Shell Game accuracy with text actions.
(b) Shell Game accuracy with image actions.
Figure 4: In the text action setting, the reasoning models Claude 3.7 Sonnet Thinking and Gemini 2.5 Pro achieve the highest performance over long action sequences. In the image action setting the accuracy of all models except Gemini 2.5 Pro decreases to random guessing by 2020 actions. Bars show 95% confidence intervals.
@@ 4 Experiments
4.1 Tracking in Text Outperforms Images
We evaluate the difference in accuracy when tracking entity states from text and image actions in the zero-shot, few-shot, chain-of-thought, and reasoning settings and report 95% confidence intervals (CI). For the few-shot experiments, $N=5$ examples are selected at random from the training set and prepended to the test example. To evaluate the effect of chain-of-thought reasoning, we prompt the model as in the zero-shot case but append “think step by step before producing a final answer.” In both domains, we evaluate on a set of $500$ games selected at random from the test set, each with a sequence length of ten actions, and $100$ games for reasoning models.
Chess
Table 1: Entity tracking accuracy (95% CI) in Chess for text-only and image-only actions. In the Few‐Shot setting $N=5$ in-context examples are used. Methods with explicit reasoning perform best. The baseline is predicting the Game Start board configuration.
Model Chess
Text Image
@@ 4.1 Tracking in Text Outperforms Images
GPT-4o 91.6 ± 0.3 76.0 ± 0.5
GPT-4o-mini 75.6 ± 0.5 55.3 ± 0.5
GPT-4.1 85.6 ± 0.4 67.7 ± 0.5
GPT-4.1-mini 82.5 ± 0.4 77.5 ± 0.5
GPT-4.1-nano 77.5 ± 0.5 73.3 ± 0.5
Gemini-2.5-Flash 91.0 ± 0.3 66.9 ± 0.5
LLaMA-4 Maverick 86.7 ± 0.4 51.1 ± 1.2
MiniMax-VL-01 85.9 ± 0.4 73.4 ± 0.5
Claude 3.7 Sonnet 96.1 ± 0.2 70.2 ± 0.5
@@
GPT-4o 94.3 ± 0.2 77.6 ± 0.5
GPT-4o-mini 74.5 ± 0.5 74.2 ± 0.5
GPT-4.1 86.6 ± 0.4 64.9 ± 0.5
GPT-4.1-mini 81.2 ± 0.4 76.9 ± 0.5
GPT-4.1-nano 74.5 ± 0.5 74.6 ± 0.5
Gemini-2.5-Flash 91.3 ± 0.3 72.0 ± 0.5
LLaMA-4 Maverick 85.8 ± 0.4 48.1 ± 0.6
MiniMax-VL-01 88.0 ± 0.8 40.8 ± 1.2
Claude 3.7 Sonnet 99.2 ± 0.1 77.7 ± 0.5
@@
GPT-4o 94.9 ± 0.2 67.5 ± 0.5
GPT-4o-mini 68.8 ± 0.5 43.1 ± 0.5
GPT-4.1 98.1 ± 0.1 75.3 ± 0.5
GPT-4.1-mini 86.4 ± 0.4 77.6 ± 0.5
GPT-4.1-nano 49.0 ± 0.6 34.6 ± 0.5
Gemini-2.5-Flash 77.0 ± 1.0 44.6 ± 1.2
LLaMA-4 Maverick 82.4 ± 0.4 61.9 ± 0.5
MiniMax-VL-01 62.3 ± 0.5 32.8 ± 0.5
Claude 3.7 Sonnet 99.5 ± 0.1 96.2 ± 0.2
@@
o1 98.2 ± 0.3 83.5 ± 0.9
o3 99.9 ± 0.1 45.5 ± 1.2
o4-mini 84.2 ± 0.9 78.0 ± 1.0
Gemini-2.5-Pro 77.4 ± 1.0 76.8 ± 1.0
Gemini-2.5-Flash 40.2 ± 1.2 17.4 ± 0.9
Claude 3.7 Sonnet 99.8 ± 0.1 96.0 ± 0.5
Claude 4.5 Opus 99.9 $\pm$ 0.1 99.5 $\pm$ 0.2
GPT-5.2 98.0 $\pm$ 0.4 99.3 $\pm$ 0.2
In the Chess domain, we report the per-square accuracy of the predicted board, that is the ratio of correctly predicted pieces (or absence of a piece) to the total number of board tiles. The ‘Game Start’ baseline is the accuracy of predicting the initial board configuration. After only ten actions, much of the board remains unchanged, so this is a strong baseline.
Table 1 reports the accuracy for text and image actions in the chess domain. Performance in the text modality is significantly better than in the image modality. In the zero-shot setting, the best-performing text model Claude 3.7 Sonnet, achieves an impressive $96.1\%\pm 0.2\%$ accuracy, while its image-based counterpart drops sharply to $70.2\%\pm 0.5\%$. A similar trend holds across models, indicating that current models can perform entity tracking in text but fail to integrate visual updates as effectively.
Both in-context learning with few-shot, chain-of-thought, and reasoning lead to performance improvements. But the text tracking accuracy exceeds the image accuracy.
Shell Game
Table 2: Entity tracking accuracy (95% CI) in Shell Game for text-only and image-only actions. In the Few‐Shot setting $N=5$ in-context examples are used. Methods with explicit reasoning perform best.
Model Shell
Text Image
@@
GPT-4o 33.0 ± 4.1 32.2 ± 4.1
GPT-4o-mini 33.6 ± 4.1 32.4 ± 4.1
GPT-4.1 32.2 ± 4.1 36.4 ± 4.2
GPT-4.1-mini 31.4 ± 4.1 30.8 ± 4.0
GPT-4.1-nano 31.2 ± 4.0 30.8 ± 4.0
Gemini-2.5-Flash 35.0 ± 4.2 37.0 ± 4.2
LLaMA-4 Maverick 30.8 ± 4.0 33.2 ± 4.1
MiniMax-VL-01 30.8 ± 4.0 31.2 ± 4.0
Claude 3.7 Sonnet 35.4 ± 4.2 37.8 ± 4.2
@@
GPT-4o 33.6 ± 4.1 32.0 ± 4.1
GPT-4o-mini 34.8 ± 4.2 32.2 ± 4.1
GPT-4.1 34.6 ± 4.2 33.0 ± 4.1
GPT-4.1-mini 36.4 ± 4.2 32.8 ± 4.1
GPT-4.1-nano 36.6 ± 4.2 36.4 ± 4.2
Gemini-2.5-Flash 31.4 ± 4.1 35.2 ± 4.2
LLaMA-4 Maverick 35.4 ± 4.2 35.2 ± 4.2
MiniMax-VL-01 36.4 ± 4.2 32.0 ± 4.1
Claude 3.7 Sonnet 32.4 ± 4.1 36.0 ± 4.2
@@
GPT-4o 99.0 ± 0.9 35.8 ± 4.2
GPT-4o-mini 61.0 ± 4.3 30.4 ± 4.0
GPT-4.1 99.8 ± 0.5 37.6 ± 4.2
GPT-4.1-mini 100.0 $\pm$ 0.4 72.0 ± 3.9
GPT-4.1-nano 61.8 ± 4.2 32.2 ± 4.1
Gemini-2.5-Flash 94.0 ± 4.8 34.0 ± 9.1
LLaMA-4 Maverick 77.0 ± 3.7 34.6 ± 4.2
MiniMax-VL-01 77.4 ± 3.7 34.4 ± 4.2
Claude 3.7 Sonnet 100.0 $\pm$ 0.4 77.4 ± 3.7
@@
o1 100.0 $\pm$ 1.9 92.6 ± 2.3
o3 100.0 $\pm$ 1.9 63.0 ± 9.3
o4-mini 100.0 $\pm$ 1.9 33.0 ± 4.1
Gemini-2.5-Pro 100.0 $\pm$ 1.9 100.0 $\pm$ 1.9
Gemini-2.5-Flash 100.0 $\pm$ 1.9 42.0 ± 4.1
Claude 3.7 Sonnet 100.0 $\pm$ 1.9 87.0 ± 6.6
Claude 4.5 Opus 100.0 $\pm$ 1.9 100.0 $\pm$ 1.9
GPT-5.2 100.0 $\pm$ 1.9 99.0 $\pm$ 2.6
@@
The final state is a single number $n\in\{1,2,3\}$ that gives the position of the ball, and we measure the accuracy of predicting it. The naive baseline picks a position uniformly at random.
@@
Table 2 reports the accuracy for text and image actions for Shell Game. The results follow a pattern similar to that of Chess, with text-based tracking outperforming image-based tracking. However, unlike in Chess, the performance in the zero-shot and few-shot settings are close to random.